### v2

# TC-GS: A Faster Gaussian Splatting Module Utilizing Tensor Cores

Abstract.
3D Gaussian Splatting (3DGS) renders pixels by rasterizing Gaussian primitives, where conditional alpha-blending dominates the computational cost in the rendering pipeline. This paper proposes TC-GS, an algorithm-independ-ent universal module that expands the applicability of Tensor Core (TCU) for 3DGS, leading to substantial speedups and seamless integration into existing 3DGS optimization frameworks. The key innovation lies in mapping alpha computation to matrix multiplication, fully utilizing otherwise idle TCUs in existing 3DGS implementations. TC-GS provides plug-and-play acceleration for existing top-tier acceleration algorithms and integrates seamlessly with rendering pipeline designs, such as Gaussian compression and redundancy elimination algorithms. Additionally, we introduce a global-to-local coordinate transformation to mitigate rounding errors from quadratic terms of pixel coordinates caused by Tensor Core half-precision computation. Extensive experiments demonstrate that our method maintains rendering quality while providing an additional 2.18× speedup over existing Gaussian acceleration algorithms, thereby achieving a total acceleration of up to 5.6×.
Gaussian Splatting, Tensor Cores, Real-Time Rendering
submissionid: 2424journalyear: 2025copyright: ccconference: arXiv Preprint Version; October 11th; 2025ccs: Computing methodologies Rasterizationccs: Computing methodologies Parallel algorithms
Figure 1. TC-GS: a Tensor Core-based acceleration module for 3D Gaussian Splatting that can be effortlessly applied to existing 3DGS rendering pipelines. Integrated into 3DGS and its variants, it achieves an extra $2\times$ speedup while preserving rendering quality.:
1. Introduction
3D Gaussian Splatting (3DGS) represents a significant advancement in neural rendering, employing Gaussian primitives to achieve NeRF-comparable quality while reducing optimization time to tens of minutes (Mallick et al., 2024; Durvasula et al., 2023). However, 3DGS rendering speed is heavily constrained by extensive model parameters and inefficient rendering pipelines(Lu et al., 2024; Ren et al., 2024; Jiang et al., 2024). Further acceleration remains essential for deployment on resource-constrained devices, real-time large-scale scene rendering systems, and low-latency edge computing scenarios.
Despite numerous efforts to enhance computational efficiency (Wang et al., 2024; Feng et al., 2024; Hanson et al., 2024; Huang et al., 2025), they fail to tap into the most powerful computing units in modern GPUs. Moreover, existing acceleration modules are tightly coupled with core components, preventing modular reuse and inter-method adaptation. This coupling further limits the practical deployment and scalability of 3DGS-based solutions.
Recent research (Wang et al., 2024; Feng et al., 2024; Hanson et al., 2024; Huang et al., 2025) has identified the conditional alpha-blending process as the primary pipeline bottleneck. This process, which composites pixel colors with depth-sorted splatted Gaussian primitives, comprises three steps: (1) Alpha computation: computing the alpha value (opacity) of each splatted Gaussian fragment on the pixel; (2) Culling: eliminating fragments with low alpha values; and (3) Blending: compositing remaining fragments into final pixel colors.
Many approaches (Ye et al., 2025; Wang et al., 2024; Fan et al., 2024) employ early culling strategies to reduce the range of splatted Gaussians, thus minimizing redundant alpha-blending computations. Although they eliminate redundant computation, these methods require corresponding pipeline designs to mitigate load imbalance issues introduced by precise intersection (Feng et al., 2024) calculations or sacrifice certain precision(Wang et al., 2024). Additionally, we observe that over 80% of Gaussian-pixel pairs that require culling remain in the pipeline. Consequently, the first two steps are the dominant part of the alpha-blending process.
Unlike previous efforts to remove redundant computing (Wang et al., 2024; Feng et al., 2024; Hanson et al., 2024; Huang et al., 2025), we focus on utilizing Tensor Core Units (TCUs) to accelerate alpha computation, which remains an unexplored area. In modern GPUs, TCUs are deployed to perform matrix multiply accumulate (MMA), i.e., $D=A\times B+C$ operations within a clock cycle, achieving high throughput on general matrix multiplication (GEMM) based computations including MLPs, CNNs, and transformers (Vaswani et al., 2017). However, due to their specialized computation pattern, TCUs struggle to be applied to non-GEMM computations, leaving them idle in 3DGS.
To tackle this problem, this paper proposes TC-GS, a hardware-aligned redesign of alpha computation that unlocks Tensor Core acceleration for all 3DGS algorithms. TC-GS formulates a mapping from pixels and Gaussians into two matrices, where their matrix product yields logarithmic alpha values in batches. Implemented as a standalone module, TC-GS not only significantly improves rendering efficiency, but also enables seamless integration with existing 3DGS acceleration kernels, further enhancing overall performance.
In summary, the paper proposes the following contributions:
  • A Detailed runtime analysis of the rendering pipeline showing that alpha computation in conditional alpha-blending is the performance bottleneck in 3DGS inference rendering.
  • The first use of Tensor Cores to accelerate alpha-blending, expanding the Tensor Core applicability beyond GEMM-based computations.
  • A global-to-local coordinate transformation (G2L, Section 4.3) that reduces the absolute magnitudes of quadratic coordinate terms, which are sensitive to Tensor Core half-precision computation, enabling lossless rendering quality while maintaining full acceleration benefits.
2. Related Works
2.1. Fast 3D Gaussian Splatting
3DGS (Kerbl et al., 2023) has emerged as a leading real-time rendering technique with photo-realistic visual quality, finding wide applications diverse domains including autonomous driving (Zhou et al., 2024), robotics (Zhu et al., 2024), and virtual reality (Jiang et al., 2024). Current acceleration research explores two primary optimization strategies to enhance 3DGS performance.
The first strategy modifies the algorithm itself to achieve better efficiency (Ye et al., 2025; Mallick et al., 2024; Lee et al., 2024). CompactGS (Lee et al., 2024), Mini-splatting (Fang and Wang, 2024), PUP 3D-GS (Hanson et al., 2025), and gsplat (Ye et al., 2025) introduce more efficient radiance field representations and adaptive resolution primitive pruning algorithms, while Taming-3DGS (Mallick et al., 2024) presents mathematically equivalent but computationally efficient solutions for gradient computation and attribute updates, substantially accelerating training speed. Stochastic rasterization for sorting-free 3DGS (Kheradmand et al., 2025) is an another approach of improving the algorithm itself.
The second category focuses on improving computational efficiency of the original algorithm through intelligent task scheduling and CUDA kernel optimization (Feng et al., 2024; Hanson et al., 2024; Wang et al., 2024; Ye et al., 2025; Gui et al., 2024; Durvasula et al., 2023). Balanced-3DGS (Gui et al., 2024) optimizes the forward computation of rendering CUDA kernels within warps while minimizing thread idle time and maximizing resource utilization by evenly distributing tasks across computational blocks. Several works, including AdRGaussian (Wang et al., 2024), FlashGS (Feng et al., 2024), and SpeedySplat (Hanson et al., 2024), aim to reduce the number of pixels processed per Gaussian by designing more precise Gaussian-tile bounding boxes, with FlashGS (Feng et al., 2024) and SpeedySplat (Hanson et al., 2024) further computing exact Gaussian-tile intersections. DISTWAR (Durvasula et al., 2023) enhances processing efficiency through sophisticated thread management strategies, fully leveraging warp-level reductions in SM sub-cores and intelligently distributing atomic computations between SM and L2 atomic units based on memory access patterns. While these methods significantly accelerate training and rendering speed, tight coupling between algorithmic modules hinders transferability and reusability. Moreover, they fail to leverage modern hardware acceleration features such as Tensor Cores, further limiting the practical deployment and scalability of 3DGS solutions.
2.2. Tensor Cores for Non-GEMM Based Computations
[…]
Despite this theoretical advantage, it is not a trival task to exploit Tensor Cores in arbitrary applications since Tensor Cores only support the matrix–multiply–accumulate (MMA) instruction rather than the full range of CUDA instructions set. This restriction requires algorithms to be reformulated in terms of small matrix operations, but the reward is exceptionally fast execution. Consequently, the strict requirements yet remarkable potential speed of Tensor Core operations have attracted researchers to explore adapting non-machine learning algorithms to GPU Tensor Cores, leveraging their performance potential. Previous works have successfully accelerated primitives such as reduction, scan and breadth-first search (BFS) using Tensor Cores (Dakkak et al., 2019a, b; Niu and Casas, 2025). Recently, ConvStencil (Liu et al., 2022) has managed to adapt stencil computation to TCUs, while other works have explored TCU-based implementations in applications such as HighQR (Leng et al., 2024) and TCUDB (Hu et al., 2022). However, to our best knowledge, no prior work has explored how to exploit Tensor Cores directly within a 3D Gaussian Splatting rendering pipeline.
Table 1. Tensor Core & CUDA Core Comparision among various GPUs. * stands for using new sparse feature
GPU Peak FP16 Peak FP16 Tensor Core
H100 (PCIe) 96 TFLOPS 800 TFLOPS — 1600 TFLOPS*
A100 78 TFLOPS 312 TFLOPS — 624 TFLOPS*
V100 31.4 TFLOPS 125 TFLOPS
RTX 4090 82.6 TFLOPS 330.3 TFLOPS — 660.6 TFLOPS*
3. Preliminaries and Observations
3.1. 3DGS Rendering Pipeline
In this paper, we focus on the official implementation of 3DGS, which is the mainstream pipeline widely used in most state-of-the-art acceleration methods.
A 3DGS model consists of a set of 3D Gaussians primitives $\mathcal{G}=\{G_{1},G_{2},\cdots,G_{P}\}$. Each Gaussian $G:\mathbb{R}^{3}\to\mathbb{R}$ is parameterized by its mean $\mu\in\mathbb{R}^{3}$, semi-definite covariance $\Sigma\in\mathbb{R}^{3\times 3}$, opacity $o\in(0,1]$, and color $c\in[0,1]^{3}$ and maps any 3D position $x\in\mathbb{R}^{3}$ to its density:
(1) $G(x)=o\mathrm{e}^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)}.$
The inference procedure of 3DGS can be divided into the following 3 stages:
Preprocess
The preprocess stage projects 3D Gaussian $G$ into 2D planes in parallel according to the viewing transform $W$ of a given camera:
(2) $\mu^{\prime}=W\mu\in\mathbb{R}^{2},\quad\Sigma^{\prime}=JW\Sigma W^{T}J^{T}\in\mathbb{R}^{2\times 2},$
where $J=\frac{\partial{\mu^{\prime}}}{\partial\mu}$ is the Jacobian of projection $W$. $\mu^{\prime}=(\mu_{x}^{\prime},\mu_{y}^{\prime})$ and $\Sigma^{\prime}$ is the mean and the covariance of projected Gaussian $G^{\prime}$. The complexity of the preprocessing stage is $\mathcal{E}_{p}=O(P)$, where $P$ is the number of Gaussians.
Sorting
The screen space $\mathcal{S}$ are divided into disjoint tiles $\mathcal{T}$ with equal sizes. For each projected Gaussian, this stage identifies all the tiles it covers. Then, for each tile $\mathcal{T}\in\mathcal{S}$, the overlapped Gaussians are sorted by depth. A tile $\mathcal{T}$ coupled with a covering Gaussian $G^{\prime}$ is denoted as a splat $(\mathcal{T},G^{\prime})$. The total number of splats given by
(3) $N=\sum_{j=1}^{P}\sum_{\mathcal{T}\in\mathcal{S}}(\mathcal{T}\cap G^{\prime}_{j}\neq\emptyset)=\sum_{\mathcal{T}\in\mathcal{S}}n(\mathcal{T}).$
where $\mathbf{1}$ is the indicator function that checks whether the Gaussian covers the tile. The process of determining coverage depends on the specific algorithm used. The complexity of the sorting stage is $\mathcal{E}_{s}=O(sN)$, where $s=64$ is the coefficient in radix-sort.
Input: Tile $\mathcal{T}$, Sorted Projected Gaussians $\mathcal{G^{\prime}(\mathcal{T})}=(G^{\prime}_{1},\cdots,G^{\prime}_{n(\mathcal{T})})$
Output: Pixel colors $C_{\mathcal{T}}=(C_{1},\cdots,C_{|\mathcal{T}|})$
foreach pixel $p\in\mathcal{T}$ in parallel do
   $T\leftarrow 1$ ;
   // Initialize transmissivity
   $C\leftarrow(0,0,0)$ ;
   // Initialize RGB
   foreach Gaussian $G^{\prime}\in\mathcal{G^{\prime}(T)}$ do
     $\alpha\leftarrow o\exp\left(-\frac{1}{2}(\mu^{\prime}-p)^{T}\Sigma^{\prime-1}(\mu^{\prime}-p)\right)$
    if $\alpha<\frac{1}{255}$ then
       continue ;
       // Cull invisible Gaussians
       end if
      if $T-\alpha{T}<0.0001$ then
         break ;
         // Terminate blending
         end if
        $C\leftarrow C+c\alpha T$ ;
         // Composite colors
         $T\leftarrow T-\alpha T$ ;
         // Update transmissivity
         end foreach
         end foreach
ALGORITHM 1 Conditional Alpha-blending
Conditional Alpha-blending
This stage renders the final color of each pixel in a tile $\mathcal{T}=(p_{1},p_{2},\cdots,p_{|\mathcal{T}|})$ in parallel as Algorithm 1. Initially, for each projected overlapping Gaussian, the renderer will perform alpha computation to get the local opacity $\alpha$ at each pixel $p=(p_{x},p_{y})\in\mathbb{R}^{2}$:
(4) $\alpha_{i,j}=o_{j}\exp{\left(-\frac{1}{2}(\mu^{\prime}_{j}-p_{i})^{T}(\Sigma^{\prime}_{j})^{-1}(\mu^{\prime}_{j}-p_{i})\right)}.$
Next, the renderer performs culling of Gaussians at each pixel if the local opacity is below a threshold:
(5) $\alpha_{i,j}<\dfrac{1}{255},$
The culling process is only a conditional judgment.
Otherwise, the surviving Gaussians will perform blending and contribute the pixel’s color:
(6) $C_{i}=\sum_{j=1}^{n(\mathcal{T})}c_{j}\alpha_{i,j}T_{i,j};\quad T_{i,j}=\prod_{k=1}^{j-1}(1-\alpha_{i,k}).$
where $T_{i,j}$ denotes the transmissivity and $n(\mathcal{T})$ is the number of splats of the tile $\mathcal{T}$. If $T_{i,j}<0.0001$, the renderer will terminate the blending process for that pixel and output the color. A pixel $p$ rendered by a Gaussian $G$ is referred to as a fragment, $(p,G)$. The complexity of the alpha-blending stage is $\mathcal{E}_{a}=O(|\mathcal{T}|N)$.
Thus, the total complexity of the Gaussian rasterization pipeline can be summarized as $\mathcal{E}=\mathcal{E}_{p}+\mathcal{E}_{s}+\mathcal{E}_{a}$, where $\mathcal{E}_{p},\mathcal{E}_{s},\mathcal{E}_{a}$ represent the complexities of preprocess, sorting, and alpha-blending stages, respectively.
3.2. Observation and Motivation
To better understand the bottlenecks in current rendering methods, we conduct a detailed time breakdown analysis of the three stages mentioned above. Specifically, we profile 3DGS, AdR-Gaussian, and Speedy-Splat on an NVIDIA A800 using multiple datasets. While recent methods incorporate early culling strategies with more precise intersection checks to reduce the number of splats $N$, thereby decreasing the time spent on sorting and blending, the alpha blending stage still remains the dominant contributor to the overall rendering time, as shown in the bottom-left corner of Figure 3.
Motivated by this observation, we further analyze the core kernel renderCUDA, which implements the alpha blending stage, in order to uncover deeper inefficiencies and guide targeted improvements.
Figure 2. a) Three types of fragments on a pixel. b) If the Gaussian only covers a small portion of tile, a large amount culled fragments are generated.
The alpha-blending stage will blend all fragments into pixel colors. These fragments are categorized into 3 types as Figure 2 shows:
  • Culled fragments: Fragments that are discarded because $\alpha<\frac{1}{255}$. As shown in Figure 2, the second fragment is culled.
  • Blended fragments: Fragments that contribute to the final pixel colors. As shown in Figure 2, the first, third, and fourth fragments are blended.
  • Skipped fragments: Fragments that are ignored when $T<0.0001$, causing the renderer to terminate the blending process for the corresponding pixel. As shown in Figure 2, the fifth fragment and the ones that follow are skipped. There is no computation on skipped fragments.
Table 2. The alpha computation and culling steps are run on both blended and culled fragments, while the blending step is only performed on blended fragments. There is no computation on skipped fragments.
Fragments Alpha Computation Culling Blending
Blended
Culled
Skipped
To model the computation for each type of fragments, we denote $k_{\alpha},k_{\text{cull}},k_{\text{blend}}$ as the average computation cost for alpha computation, culling, and blending. Similarly, let $f_{\text{blend}},f_{\text{cull}},f_{\text{skip}}$ represent the number of blended, culled and skipped fragments. According to Table 2, the total computation amount is:
(7) $\mathcal{C}=k_{\text{blend}}f_{\text{blend}}+(k_{\text{cull}}+k_{\alpha})(f_{\text{blend}}+f_{\text{cull}}).$
To identify the bottleneck within the alpha-blending stage, we collect and categorize the generated fragments on various scenes using 3DGS, AdR-Gaussian and Speedy-Splat. As shown on the right side of Figure 3, culled and skipped fragments account for a large proportion. Since skipped fragments do not incur any computational cost, we can preliminarily conclude that the bottleneck lies in the alpha computation and culling steps.
Figure 3. Analysis of 3DGS rendering bottlenecks: The left side shows the time distribution of preprocessing, sorting, and alpha-blending for 3DGS, AdR-Gaussian, and Speedy-Splat, with alpha-blending dominating. The right side details alpha computation, culling, and blending, identifying culled fragments as the primary bottleneck due to redundant alpha computations, while skipped fragments incur no cost.
4. Method
Figure 4. Design of Frag2Mat: the alpha computation is reformulated as matrix multiplication to fully leverage Tensor Cores for accelerating alpha calculation.
This section introduces how TC-GS utilizes the Tensor Cores to accelerate the alpha-computation and culling process, which are the bottlenecks of rendering. The fundamental components of TC-GS include EarlyCull for culling optimization, Frag2Mat for matrix transformation and G2L for coordinate transformation. The detailed design is shown in Figure 4.
4.1. EarlyCull: Reducing Exponential Computation
First, TC-GS adopts EarlyCull, an early-pruning strategy: it identifies invisible Gaussians that can be culled before the alpha-computation stage, greatly reducing the number of exponential instructions exp executed and improving throughput.
We begin by expressing Equation (4) as the following form:
(8) $\alpha_{i,j}=\mathrm{e}^{\beta_{i,j}},$
where the exponent $\beta_{i,j}$ satisfies
(9) $\beta_{i,j}=\ln(o_{j})-\frac{1}{2}(\mu^{\prime}_{j}-p_{i})^{T}(\Sigma^{\prime}_{j})^{-1}(\mu^{\prime}_{j}-p_{i}).$
Then the culling condition, Equation (5), can be converted into
(10) $\beta_{i,j}<-\ln(255).$
EarlyCull allows us to filter out non-contributing Gaussians in advance, reducing the amount of expensive exp instructions issued by alpha computation. Since prior experiments show that culled fragments account for a large portion of the workload, this approach yields a substantial speed-up for the entire pipeline.
4.2. Frag2Mat: Fitting the MMA Operation
While Tensor Cores specialize in matrix multiply-accumulate (MMA) operations, the original alpha-computation involves per-pixel-per-Gaussian evaluations that exhibit non-coalesced memory access patterns and fragmented arithmetic operations – characteristics mismatched with matrix-oriented hardware architectures. To align with Tensor Cores’ computational paradigm and fully exploit their high throughput, we restructure the alpha-computation into batched matrix operations through algebraic reformulation.
We propose Frag2Mat, a computational restructuring that transforms per-pixel-per-Gaussian exponent calculations into unified matrix operations. By expressing the exponent as a linear combination of pixel coordinates, Gaussian parameters and their quadratic terms, we decouple multivariate dependencies into two independent components: a pixel vector encoding coordinate polynomials for its pixel, and a Gaussian vector composed of parameters from its Gaussian. The final exponent values are batch-calculated through standard matrix multiplication of two matrices formed by stacking pixel vectors and Gaussian vectors respectively. LC The core idea of TC-GS is decoupling the computation of exponents $\beta_{i,j}$ into two separated terms: one that depends only on the pixel index and another that depends only on the Gaussian. Notably, we observe that the exponent can be expressed as a quadratic function of the pixel $p$. Based on this observation, TC-GS transforms Equation (9 into the following standard from:
(11) $\beta=v_{0}+v_{1}p_{x}+v_{2}p_{y}+v_{3}p_{x}^{2}+v_{4}p_{x}p_{y}+v_{5}p_{y}^{2},$
where
(12) $v_{0}$ $=\ln(o)-\frac{1}{2}\mu^{\prime T}\Sigma^{\prime-1}\mu^{\prime},$
$v_{1}$ $=\sigma_{11}\mu^{\prime}_{x}+\sigma_{12}\mu^{\prime}_{y},\quad v_{2}=\sigma_{12}\mu^{\prime}_{x}+\sigma_{22}\mu^{\prime}_{y},$
$v_{3}$ $=-\dfrac{1}{2}\sigma_{11},\quad v_{4}=-\sigma_{12},\quad v_{5}=-\dfrac{1}{2}\sigma_{22}$
with the notation that symmetric $\Sigma^{\prime-1}=\begin{bmatrix}\sigma_{11}&\sigma_{12}\\ \sigma_{21}&\sigma_{22}\end{bmatrix}$ with $\sigma_{12}=\sigma_{21}$.
The Equation (11) is the expanded dot product of two vectors:
(13) $\beta_{i,j}=\mathbf{u}_{i}^{T}\mathbf{v}_{j},$
where
(14) $\mathbf{u}=\begin{bmatrix}1&p_{x}&p_{y}&p_{x}^{2}&p_{x}p_{y}&p_{y}^{2}\end{bmatrix}^{T}$
and
(15) $\mathbf{v}=\begin{bmatrix}v_{0}&v_{1}&v_{2}&v_{3}&v_{4}&v_{5}\end{bmatrix}^{T}.$
Here $\mathbf{u}$, the pixel vector solely dependent on pixel properties, which will be computed only once for each pixel. Similarly, each Gaussian vector $\mathbf{v}$ can be only computed once. Therefore, the expression of the exponent is decoupled into the dot product of two individual vectors. To align the computation to the Tensor Cores, TC-GS stacks the vectors into two independent matrices:
(16) $\mathbf{U}=\begin{bmatrix}\mathbf{u}_{1}&\mathbf{u}_{2}&\cdots&\mathbf{u}_{m}\end{bmatrix}\in\mathbb{R}^{6\times m},$
(17) $\mathbf{V}=\begin{bmatrix}\mathbf{v}_{1}&\mathbf{v}_{2}&\cdots&\mathbf{v}_{j}\end{bmatrix}\in\mathbb{R}^{6\times n},$
where $m$ is the the number of pixels within a tile, and $n$ is the number of Gaussians whose projection is overlapped with the tile. Then TC-GS computes the exponent matrix $\mathbf{B}\in\mathbb{R}^{m\times n}$ as matrix multiplication:
(18) $\mathbf{B}=\begin{bmatrix}\beta_{1,1}&\cdots&\beta_{1,n}\\ \vdots&\ddots&\vdots\\ \beta_{m,1}&\cdots&\beta_{m,n}\\ \end{bmatrix}=\begin{bmatrix}\mathbf{u}_{1}^{T}\mathbf{v}_{1}&\cdots&\mathbf{u}_{1}^{T}\mathbf{v}_{n}\\ \vdots&\ddots&\vdots\\ \mathbf{u}_{m}^{T}\mathbf{v}_{1}&\cdots&\mathbf{u}_{m}^{T}\mathbf{v}_{n}\\ \end{bmatrix}=\mathbf{U}^{T}\mathbf{V}.$
Therefore, we can leverage Tensor Cores to compute $\beta$ for $m$ pixels and $n$ Gaussians simultaneously. As Tensor Cores natively support half-precision matrix operations, additional design is required to ensure numerical stability throughout the alpha computation.
4.3. G2L: Fitting the Half Precision
Although we adopted Tensor Cores for alpha computation and culling. The challenges remain as Tensor Cores do not support common single-precision floating-point operations. Instead, they only take FP16 or TF32 as the input of the MMA operation, resulting in a higher machine epsilon of $\varepsilon_{H}=9.77\times 10^{-4}$ compared to FP32’s $\varepsilon_{F}=1.19\times 10^{-7}$. The rounding error of a floating number $r$ is $|r|\varepsilon_{H}$. Therefore, the error of computation on Tensor Cores is more sensitive at the absolute value of the input data.
We notice that the range of the pixel $p$ is $0,w]\times[0,h]$, where $w,h$ are the width and height of the screen space, respectively. When loading the pixel vector $\mathbf{u}$ in Equation (14) into Tensor Cores, its absolute values of quadratic terms $p_{x}^{2}$, $p_{x}p_{y}$ and $p_{y}^{2}$ can exceed a million, contributing catastrophic rounding errors to $\beta$. Moreover, overflow will occur if the FP16 range $6.55\times 10^{5},6.55\times 10^{5}]$ is used.
Figure 5. The quadratic term of pixel coordinates significantly contributes the rounding error, resulting in blurry images. Local coordinates constrain the value of $\Delta p$ into $8,8]^{2}$, which can reduce the upper bound of the rounding error.
To tackle this problem, TC-GS has designed G2L, a coordinate mapping, which significantly reduces the absolute value of terms in pixel vectors. Since all pixels $p$ within a tile $\mathcal{T}$ will be rendered in parallel, G2L can map the coordinates of pixels into the local coordinate:
(19) $\Delta p=p-p(\mathcal{T}),$
Since each tile contains $16\times 16$ pixels, then the range of pixel in LC is $\Delta p\in[-8,8]^{2}$, and the maximum element of $\mathbf{u}_{\Delta}$ will not exceed $64$. This conversion does not only avoid the overflow in FP16, but also significantly reduces the rounding error.
Respectively, the means of Gaussians in the local coordinate is
(20) $\Delta\mu^{\prime}=\mu^{\prime}-p(\mathcal{T}).$
Notice that the relative position between $\mu^{\prime}$ and $p$ stays invariant:
(21) $\Delta\mu^{\prime}-\Delta p=\mu^{\prime}-p.$
Therefore, Frag2Mat can be also performed on local coordinates.
As the proof in the appendices goes, G2L reduces the quadratic rounding error $O(h^{2}+hw+w^{2})\varepsilon_{H}$ into linear $O(h+w)\varepsilon_{H}$, and achieve better rendering quality in FP16 as Section 5.3 demonstrates.
5. Experiment
5.1. Experiment Setup
Table 3. Comparison With/Without TC-GS Module Across Datasets.
Method Tanks & Temple Deep-Blending Mip-NeRF360
[…]
3DGS 161.548 23.687 0.851 0.169 131.373 29.803 0.907 0.238 128.88 26.546 0.785 0.25
[…]
FlashGS 326.835 23.706 0.851 0.169 563.590 29.789 0.907 0.239 399.772 26.551 0.786 0.251
[…]
Speedy-Splat 268.602 23.687 0.852 0.17 315.822 29.803 0.907 0.238 264.23 26.546 0.785 0.25
[…]
AdR-Gaussian 278.65 23.635 0.852 0.17 324.18 29.814 0.907 0.238 261.339 26.52 0.785 0.25
[…]
Implementation
Current computation patterns which Tensor Cores support do not include length-6 MMA operations. A simple way to align the required length-8 MMA operation is zero padding. In our implementation, we replace $v_{0}$ in $\mathbf{v}$ with $\frac{v_{0}}{3}$ and duplicate it 3 times. The constant term $1$ in $\mathbf{u}$ are duplicated respectively. This approach can reduce the rounding error by reducing the absolute value.
datasets&metrics
To prove that our method is universal and efficient, we conduct experiments on the same datasets as those used in 3DGS(Kerbl et al., 2023). Specifically, the datasets contain all scenes from Mip-NeRF360(Barron et al., 2022), two outdoor scenes (Truck & Train) from Tanks & Temple(Knapitsch et al., 2017), two indoor scenes (Drjohnson & playroom) in Deep-Blending(Hedman et al., 2018). For each scene, CUDA events are inserted at the start and end of the forward rendering procedure to measure FPS values, and the reported results represent the average metrics of the test datasets.All experiments are conducted on an NVIDIA A800 GPU (80GB SXM) equipped with 432 Tensor Cores, achieving a peak 624 TFLOPS FP16 performance.
5.2. Comparisons
TC-GS demonstrates significant performance improvements in all tested methods while keeping rendering quality. The qualitative and quantitative comparison between four baseline rendering pipelines and their results after integrating the TC-GS module are shown in Table 3.
Rendering speed
[…]
Image Quality
Table 3 also demonstrates the rendered images quality metrics across all methods. We observe negligible differences in PSNR, SSIM, and LPIPS metrics between each method’s original implementation and its TC-GS-enhanced counterpart. The minimal variations observed likely originate from hardware-level instruction set disparities and floating-point precision conversion artifacts in the rasterization pipeline.
Table 4. Comparison of alpha-blending time across methods and datasets.
Method Scene with TC-GS time(ms) original time(ms)
[…]
5.3. Ablation Study
To determine the source of the acceleration, we conducted an ablation study on EarlyCull and Frag2Mat with G2L across various renderers and datasets. As Table 5 shows, both methods can accelerate the rendering speed. EarlyCull alone achieves $1.06-1.51\times$ speedup, but it accelerates slightly when Frag2Mat is enabled. Frag2Mat alone achieves $1.97-2.24\times$ speedup, and it still accelerates significantly with EarlyCull enabled. Therefore, Frag2Mat with G2L contributes the main speedup.
Table 5. Rendering FPS across settings, renderers, and datasets. To evaluate the effect of each component in TC-GS, we conducted an ablation study on EarlyCull and Frag2Mat with G2L.
EarlyCull Frag2Mat with G2L 3DGS SpeedySplat
truck drjohnson truck drjohnson
169.834 108.123 366.457 349.673
210.110 163.967 406.434 371.605
351.965 242.993 732.008 688.217
361.839 253.770 727.647 701.254
To clearly illustrate the impact of G2L on rendering quality and performance, we ran the following comparisons shown in Table 6 and the rendering images are shown in Figure 6. The ablation experiment shows that all Tensor Core variants boosted FPS. However, direct TF32 computation already introduced noticeable image-quality degradation, and FP16 alone dropped PSNR to just 8 dB. In contrast, combining FP16 Tensor Core optimization with Local Coordinates method preserved nearly same image quality while delivering the highest FPS. This demonstrates that our G2L can simultaneously maximize performance and maintain rendering fidelity.
Table 6. Ablation study on applying TC-GS on original 3DGS.
Frag2Mat G2L Precision drjohnson truck
$uparrow PSNR\uparrow FPS\uparrow$
FP32 29.48 108.123 25.44 169.834
$\checkmark$ FP16 8.85 253.79 6.24 356.619
$\checkmark$ TF32 20.02 181.429 14.27 290.047
$\checkmark$ $\checkmark$ FP16 29.46 253.77 25.44 361.839
Figure 6. Ablation study on G2L when applying TC-GS on original 3DGS.
5.4. Discussion
5.4.1. Limitations
While our method achieves significant acceleration in rendering computations, further optimization could be attained by adopting pipeline techniques similar to those proposed in FlashGS. However, unlike FlashGS’s thread-level pipelining, a warp-level pipelining approach is required. Notably, post-optimization profiling reveals an increased proportion of preprocess overhead in the forward procedure. This preprocess phase inherently involves numerous matrix-based operations such as view transformations, providing opportunities for optimization using Tensor Cores.
5.4.2. Training
Although TC-GS is primarily designed to optimize real-time rendering, our implementation serves as a plug-and-play accelerator for both inference and training, requiring no modifications to the original 3DGS model or the training process.
6. Conclusion
In this paper, we propose TC-GS, a high-performance rendering module for accelerating 3D Gaussian Splatting. TC-GS optimizes the rendering pipeline by decomposing alpha-blending into two components, i.e., Gaussian-related and pixel-related matrices. Through reformulating alpha-blending as matrix multiplications, TC-GS
[…]
Acknowledgements.
This work is supported by the National Key R&\&D Program of China (No. 2022ZD0160201), Shanghai Artificial Intelligence Laboratory, and Shanghai Municipal Science and Technology Major Project.
Appendix A Derivation and Proof of the Upper Bound of Rounding Error on TC-GS
A.1. Floating-Point Numbers
In computing, floating-point numbers arithmetic is arithmetic on subset of real numbers, which can be represented with a finite amount of memory. A floating-point number $r$ is formed by a significand $s$, base $b$ and exponent $e$:
(22) $r=s\cdot b^{e}.$
In this paper, we only consider the binary formats, whose base $b$ is $2$.
In most modern computing systems, a floating-point number is consists of 3 parts:
  • Sign bit: Represents the sign of the number. A sign bit of 0 indicates a positive number, and $1$ indicates a negative number.
  • Exponent: Represents the scale of the number. The exponent is stored with an offset.
  • Significand: Represents the mantissa of the floating-point number. The number of bits of significand is the precision.
For example, if the significand has $m$ bits, the real value of the floating-point number is:
(23) $(-1)^{\text{sign}}(1.b_{m-1}b_{m-2}\dots b_{0})_{2}\times 2^{\text{exponent - offset}},$
where $b_{m-1}b_{m-2}\dots b_{0}$ is the significand bits. Table 7 shows 3 floating-point formats in this paper. FP32 is also known as single-precision floating-point format, and FP16 is called half-precision floating-point format.
Format Sign Exponents Significand Machine Epsilon
FP16 11bit 55bits 1010bits $\varepsilon_{H}=9.77\times 10^{-4}$
TF32 11bit 88bits 1010bits $\varepsilon_{H}=9.77\times 10^{-4}$
FP32 11bit 88bits 2323bits $\varepsilon_{F}=1.19\times 10^{-7}$
Table 7. 3 binary floating-point formats in this paper.
A.2. Machine Epsilon
Rounding is a procedure for representation of a real number $r$ $(r\neq 0)$ into a floating-point format as $R(r)$, where $R$ is the rounding function. The relative error $\varepsilon$ of the rounding procedure is defined as:
(24) $\varepsilon(r)=\dfrac{|R(r)-r|}{|r|}.$
For a fixed floating-point system and a rounding procedure, machine epsilon $\varepsilon_{\text{mach}}$ is an upper bound of relative error if no underflow or overflow occurs:
(25) $|R(r)-r|\leq\varepsilon_{\text{mach}}|r|.$
Proposition A.1.
If $\varepsilon_{\text{mach}}$ is a machine epsilon of a floating-point format, the rounded value of a real number $r$ satisfies:
(26) $R(r)=r+O(\varepsilon_{\text{mach}}).$
Proof.
Since the $\varepsilon_{\text{mach}}$ does not depend on $r$, then Inequality (25) shows $|R(r)-r|=O(\varepsilon_{\text{mach}})$. Therefore, $R(r)-r=O(\varepsilon_{\text{mach}})$.
Proposition A.2.
If a binary floating-point format has $m$ significand bits, $\varepsilon_{\text{mach}}=2^{-m}$ is its machine epsilon.
Proof.
Considering $|r|\in[2^{e},2^{e+1})$ is a real number which does not overflow or underflow in this format, it will lies between two adjacent floating-point numbers, $r_{b}$ and $r_{u}$:
(27) $r_{b}\leq|r|<r_{u}.$
Suppose $x_{b}=(1.b_{m-1}b_{m-2}\dots b_{0})_{2}\times 2^{e}$, then the adjacent $r_{u}$ is:
(28) $r_{u}=(1.b_{m-1}b_{m-2}\dots b_{0})_{2}+(0.00\cdots 1)_{2}\times 2^{e}=r_{b}+2^{e-m}.$
Since $R(r)$ is either $r_{u}$ or $r_{b}$,
(29) $|R(r)-r|\leq|r_{u}-r_{b}|=2^{e-m}=2^{e}\cdot 2^{-m}\leq 2^{-m}|r|.$
Therefore, $2^{-m}$ is a machine epsilon of the floating-point format.
From Proposition A.2, we can conclude that the machine epsilon is only depend on the number of significand bits.
In this paper, we represent the machine epsilon of TF32 and FP16 as $\varepsilon_{H}=9.77\times 10^{-4}$, and the machine epsilon of FP32 as $\varepsilon_{F}=1.19\times 10^{-7}$. Meanwhile, the rounding process of FP16 and TF32 is denoted as $R_{H}$, while the rounding process of FP32 is represented as $R_{F}$.
Furthermore, if the format is fixed, the rounding error is only depend on the absolute value $|r|$. The larger absolute value contribute the larger rounding error. The core idea of G2L in Section 4.3 is reducing the absolute value during the computing process.
A.3. Rounding Error of Dot Product with Tensor Cores
Matrix multiplication can be interpreted as the dot product of vectors at each element, as discussed in Section 4.2. Therefore, to analyze the rounding error in matrix multiplication, it suffices to examine the error in the dot product:
(30) $\beta=\mathbf{u}^{T}\mathbf{v}=\sum_{i=1}^{n}u_{i}v_{i},$
where $n$ is the length of the vectors.
The dot product can be decomposed as a multiplication step and a accumulation step. On Tensor Cores, the multiplication step are performed on half-precision format (FP16 or TF32) and the accumulation step can be executed on FP32 format.
We omit the detailed analysis on terms of small $\varepsilon_{H}^{2}$ and $\varepsilon_{F}$ and represent them as $O(\varepsilon_{H}^{2})$ and $O(\varepsilon_{F})$.
Proposition A.3.
On Tensor Cores, the absolute error of $\beta=\mathbf{u}^{T}\mathbf{v}$ satisfies
(31) $|\hat{\beta}-\beta|\leq 2\varepsilon_{H}\sum_{i=1}^{n}|u_{i}v_{i}|+O(\varepsilon_{H}^{2}+\varepsilon_{F}),$
where $\hat{\beta}$ is the computed value of $\beta$ with Tensor Cores.
Proof.
At first, we consider product of $u_{i}$ and $v_{i}$. Before multiplication, these two real number are stored in half-precision format as $R_{H}(u_{i})$ and $R_{H}(v_{i})$ satisfying
(32) $|R_{H}(u_{i})|\leq|R_{H}(u_{i})-u_{i}|+|u_{i}|\leq|u_{i}|\varepsilon_{H}+|u_{i}|=(1+\varepsilon_{H})|u_{i}|$
and
(33) $|R_{H}(v_{i})|\leq(1+\varepsilon_{H})|v_{i}|.$
Therefore,
(34) $|R_{H}(u_{i})R_{H}(v_{i})|\leq(1+\varepsilon_{H})^{2}(|v_{i}||u_{i}|)=(1+2\varepsilon_{H})|u_{i}v_{i}|+O(\epsilon_{H}^{2}).$
The value product are stored in FP32 format for next accumulation step on Tensor Cores:
(35) $|R_{F}(R_{H}(u_{i})R_{H}(v_{i}))|$ $\leq(1+\varepsilon_{F})|R_{H}(u_{i})R_{H}(v_{i})|$
$\leq(1+\varepsilon_{F})((1+2\varepsilon_{H})|u_{i}v_{i}|+O(\varepsilon_{H}^{2}))$
$\leq(1+2\varepsilon_{H})|u_{i}v_{i}|+O(\varepsilon_{H}^{2})+O(\varepsilon_{F}).$
Notice that the rounding function does not change the sign of the number. Therefore, relative error of the product is:
(36) $\varepsilon(u_{i}v_{i})$ $=\dfrac{|u_{i}v_{i}-R_{F}(R_{H}(u_{i})R_{H}(v_{i}))|}{|u_{i}v_{i}|}$
$=\dfrac{||u_{i}v_{i}|-|R_{F}(R_{H}(u_{i})R_{H}(v_{i}))||}{|u_{i}v_{i}|}$
$\leq 2\varepsilon_{H}+O(\varepsilon_{H}^{2})+O(\varepsilon_{F}).$
To simply the proof, we represent $a_{i}$ as the accumulator on Tensor Cores:
(37) $a_{i}=R_{F}(R_{H}(u_{i})R_{H}(v_{i}))=u_{i}v_{i}(1\pm\varepsilon(u_{i}v_{i}))$
We assume that the addition is performed from left to right. Since the sum of each two values are stored in FP32 as a new accumulator, the computed cumulative sum $s_{i}$ of $\{a_{i}\}_{i=1}^{n}$ is:
(38) $s_{1}=R_{F}(a_{1})=a_{1},\quad s_{i}=R_{F}(s_{i-1}+a_{i}).$
Therefore, the computed value of $\beta$ on Tensor Cores is:
(39) $\hat{\beta}=R_{F}(R_{F}(\cdots R_{F}(R_{F}(a_{1}+a_{2})+a_{3})\cdots)+a_{n})=s_{n}.$
Next, we will prove that $s_{i}=\sum_{j=1}^{i}a_{j}+O(\varepsilon_{F})$ inductively:
  1. (1)
    It is obvious that $s_{1}=a_{1}=a_{1}+O(\varepsilon_{F})$;
  2. (2)
    For a positive integer $i<n$, we assume that $s_{i}=\sum_{j=1}^{i}a_{j}+O(\varepsilon_{F})$. According to Proposition 26,
    (40) $s_{i+1}=R_{F}(s_{i}+a_{i})=s_{i}+a_{i}+O(\varepsilon_{F})=\sum_{j=1}^{i+1}a_{j}+O(\varepsilon_{F}).$
According to Equation (37),
(41) $\hat{\beta}=s_{n}=\sum_{i=1}^{n}a_{i}+O(\varepsilon_{H})=\sum_{i=1}^{n}(u_{i}v_{i}\pm\varepsilon(u_{i}v_{i}))+O(\varepsilon_{F}).$
To sum up, the absolute error of $\beta$ satisfies
(42) $|\hat{\beta}-\beta|$ $=\left|\sum_{i=1}^{n}u_{i}v_{i}(1\pm\varepsilon(u_{i}v_{i}))+O(\varepsilon_{F})-\sum_{i=1}^{n}u_{i}v_{i}\right|$
$=\left|\sum_{i=1}^{n}\pm u_{i}v_{i}\varepsilon(u_{i}v_{i})+O(\varepsilon_{F})\right|$
$=\left|\sum_{i=1}^{n}\pm u_{i}v_{i}(2\varepsilon_{H}+O(\varepsilon_{H}^{2})+O(\varepsilon_{F}))+O(\varepsilon_{F})\right|$
$=\left|\sum_{i=1}^{n}\pm 2u_{i}v_{i}\varepsilon_{H}+O(\varepsilon_{H}^{2})+O(\varepsilon_{F})\right|$
$\leq 2\varepsilon_{H}\sum_{i=1}^{n}|u_{i}v_{i}|+O(\varepsilon_{H}^{2}+\varepsilon_{F}).$
The proof of Proposition A.3 is completed here.
A.4. Absolute Error of $\beta$ on Global Coordinates
This section discusses the absolute of exponent $\beta$ on global coordinates.
The absolute error of $\beta$ is dependent on $\mathbf{v}$, and $\mathbf{v}$ is depend on the parameters of projected Gaussian $G^{\prime}$. Therefore, we discusses the range of Gaussians’ parameters at first.
We represent the projected Gaussian as $G^{\prime}=(\mu^{\prime},\Sigma^{\prime},o,c)$ with $\mu^{\prime}=(\mu^{\prime}_{x},\mu^{\prime}_{y})$ and symmetric $\Sigma^{\prime-1}=\begin{bmatrix}\sigma_{11}&\sigma_{12}\\ \sigma_{21}&\sigma_{22}\end{bmatrix}$ satisfying $\sigma_{12}=\sigma_{21}$. The alpha value at pixel $p=(p_{x},p_{y})$ is $\alpha=\mathrm{e}^{\beta}=\mathrm{e^{\mathbf{u^{T}}\mathbf{v}}}$ as Section 4.2 discussed.
The width and height of rendered image are represented as $w$ and $h$ respectively.
Means $\mu^{\prime}$:
To simplify the analysis, we only consider Gaussians whose center within the screen space. Therefore, $\mu^{\prime}\in[0,w]\times[0,h]$.
Opacity $o$:
The fragment with alpha value below $\frac{1}{255}$ will be culled. Since $\alpha\leq o$ since $\Sigma^{\prime-1}$ is semi-definite, $o\in[\frac{1}{255},1]$.
Inverse Covariance $\Sigma^{\prime-1}$:
The covariance $\Sigma^{\prime}$ identifies the shape of the ellipse on the screen space. If $\lambda_{1}$ and $\lambda_{2}$ are the two eigenvalues of $\Sigma^{\prime}$, then length of two semi-axes of the ellipse are $\lambda_{1}$ and $\lambda_{2}$.
We assume the length of semi-axis is larger than 0.5, otherwise the Gaussian will not be rendered:
(43) $\lambda_{1},\lambda_{2}\geq 0.5.$
Then the eigenvalues of the inverse covariance $\Sigma^{\prime-1}$ satisfies:
(44) $0<\lambda_{1}^{-1}+\lambda_{2}^{-1}\leq 4,\quad 0<\lambda_{1}^{-1}\lambda_{2}^{-1}\leq 4.$
Consider the characteristic polynomial of $\Sigma^{\prime-1}$:
(45) $f(\lambda^{-1})=|\lambda^{-1}I-\Sigma^{-1}|=\lambda^{-2}-(\sigma_{11}+\sigma_{22})\lambda^{-1}+\sigma_{11}\sigma_{22}-\sigma_{12}^{2}.$
These two eigenvalues satisfy $f(\lambda^{-1})=0$. Applying the Vieta’s Formulas,
(46) $0<\sigma_{11}+\sigma_{22}\leq 4,\quad 0<\sigma_{11}\sigma_{22}-\sigma_{12}^{2}\leq 4.$
Then,
(47) $\sigma_{12}^{2}<\sigma_{11}\sigma_{22}\leq\left(\frac{\sigma_{11}+\sigma_{22}}{2}\right)^{2}\leq 4.$
Therefore,
(48) $0<\sigma_{11}<4,\quad-2\leq\sigma_{12}\leq 2,\quad 0<\sigma_{22}<4.$
According to Proposition A.3, the absolute error of $\beta$ on Frag2Mat is:
(49) $|\hat{\beta}-\beta|\leq 2\varepsilon_{H}\sum_{i=0}^{5}|u_{i}v_{i}|+O(\varepsilon_{H}^{2}+\varepsilon_{F}).$
Applying the global coordinate, the coordinate of pixel satisfies $p=(p_{x},p_{y})\in[0,w]\times[0,h]$.
Then,
(50) $|u_{0}v_{0}|=\left|\ln(o)-\frac{1}{2}\sigma_{11}\mu_{x}^{2}-\sigma_{12}\mu_{x}\mu_{y}-\sigma_{22}\mu_{y}^{2}\right|=O(w^{2}+wh+h^{2}),$
(51) $|u_{1}v_{1}|=|p_{x}(\sigma_{11}\mu_{x}+\sigma_{12}\mu_{y})|=O(w^{2}+wh),$
(52) $|u_{2}v_{2}|=|p_{y}(\sigma_{12}\mu_{x}+\sigma_{22}\mu_{y})|=O(wh+h^{2}),$
(53) $|u_{3}v_{3}|=\left|p_{x}^{2}(-\frac{1}{2}\sigma_{11})\right|=O(w^{2}),$
(54) $|u_{4}v_{4}|=|p_{x}p_{y}(-\sigma_{12})|=O(wh),$
(55) $|u_{5}v_{5}|=\left|p_{y}^{2}(-\frac{1}{2}\sigma_{22})\right|=O(h^{2}).$
Therefore, the absolute error of $\beta$ satisfies quadratic
(56) $|\hat{\beta}-\beta|=O(w^{2}+wh+h^{2})\varepsilon_{H}+O(\varepsilon_{H}^{2}+\varepsilon_{F}).$
A.5. Absolute Error of $\beta$ on Local Coordinates
This section discusses the absolute of exponent $\beta$ on local coordinates.
At first, we discuss the range of Gaussians’ parameters as Section A.4. In this section, $p(\mathcal{T})$ is the center of tile $\mathcal{T}$, and the size of each tile is $16\times 16$.
Means $\Delta\mu^{\prime}$:
Since $\mu^{\prime},p(\mathcal{T})\in[0,w]\times[0,h]$, $\Delta\mu^{\prime}=\mu^{\prime}-p(\mathcal{T})\in[-w,w]\times[-h,h]$.
Opacity $o$:
The opacity is the same as Gaussians on global coordinates. $o\in[\frac{1}{255},1]$.
Inverse Covariance $\Sigma^{\prime-1}$:
. The covariance is the same as Gaussians on global coordinates: $0<\sigma_{11},\sigma_{12}<4,-2\leq\sigma_{12}\leq 2$.
All pixels on local coordinates is within a tile, then $\Delta p=(\Delta p_{x},\Delta p_{y})\in[-8,8]^{2}$.
Therefore,
(57) $|u_{1}v_{1}|=|\Delta p_{x}(\sigma_{11}\Delta\mu_{x}+\sigma_{12}\Delta\mu_{y})|=O(w+h),$
(58) $|u_{2}v_{2}|=|\Delta p_{y}(\sigma_{12}\Delta\mu_{x}+\sigma_{22}\Delta\mu_{y})|=O(w+h),$
(59) $|u_{3}v_{3}|=\left|\Delta p_{x}^{2}(-\frac{1}{2}\sigma_{11})\right|=O(1),$
(60) $|u_{4}v_{4}|=|\Delta p_{x}\Delta p_{y}(-\sigma_{12})|=O(1),$
(61) $|u_{5}v_{5}|=\left|\Delta p_{y}^{2}(-\frac{1}{2}\sigma_{22})\right|=O(1).$
According to Proposition A.3,
(62) $|\hat{\beta}-\beta|\leq(2|u_{0}v_{0}|+O(w+h))\varepsilon_{H}+O(\varepsilon_{H}^{2}+\varepsilon_{F}).$
We only care the precision of fragments which are not culled, then $\ln(255)\leq\beta<0$. Since $\beta=\sum_{i=0}^{5}u_{i}v_{i}$, then $u_{0}v_{0}=\beta-\sum_{i=1}^{5}u_{i}v_{i}$. Therefore,
(63) $|u_{0}v_{0}|\leq|\beta|+\sum_{i=1}^{5}|u_{i}v_{i}|=O(w+h).$
In conclusion, the absolute error of $\beta$ satisfies
(64) $|\hat{\beta}-\beta|\leq O(w+h)\varepsilon_{H}+O(\varepsilon_{H}^{2}+\varepsilon_{F}).$
Appendix B Notations
To clearly present the various parameters and their definitions involved in this paper, the following parameter table lists each symbol along with its corresponding meaning, facilitating readers’ understanding of the subsequent analysis and computational processes.
Notation Meaning Notation Meaning
$\mathcal{G}$ Set of all 3D Gaussian $G$ 3D Gaussian
$P$ Number of 3D Gaussians $\mu$ Mean of 3D Gaussian
$\Sigma$ Covariance of Covariance $o$ Opacity of 3D Gaussian
$c$ Color of 3D Gaussian $W$ Matrix projecting 3D position into 2D plane
$J$ Jacobian of projection $\mu^{\prime}=(\mu^{\prime}_{x},\mu^{\prime}_{y})$ Mean of projected Gaussian
$\Sigma^{\prime}=\begin{bmatrix}\sigma_{11}&\sigma_{12}\\ \sigma_{21}&\sigma_{22}\end{bmatrix}^{-1}$ Covariance of projected Gaussian $G^{\prime}$ Projected Gaussian
$\mathcal{G}^{\prime}$ Set of projected Gaussians $p=(p_{x},p_{y})$ Pixel
$\mathcal{T}$ Tile, containing pixels $\mathcal{S}$ Screen space, containing all tiles
$C$ Rendered pixel color $\alpha$ Alpha value of fragments.
$T$ Transmissivity of fragments. $n(\mathcal{T})$ Number of splats on tile $\mathcal{T}$
NN Number of splats. $\mathcal{E}$ Complexity of the pipeline in 3DGS
$w$ Width of the screen $h$ Height of the screen
$\mathcal{E}_{p}$ Complexity of preprocessing $\mathcal{E}_{s}$ Complexity of sorting
$\mathcal{E}_{a}$ Complexity of alpha-blending $k_{\alpha}$ Computation amount in alpha computation
$k_{\text{cull}}$ Computation amount in culling $k_{\text{blend}}$ Computation amount in blending
$f_{\text{cull}}$ Number of culled fragments $f_{\text{blend}}$ Number of blended (shaded) fragments.
$f_{\text{skip}}$ Number of skipped fragments $\mathcal{C}$ Total computation amount in alpha-blending.
$\mathbf{u}$ Pixel vector $\mathbf{v}$ Gaussian vector
$(p_{x},p_{y})$ Coordinates of pixel $p$ $\beta$ $\ln(\alpha)$, the exponent
$\mathbf{U}$ Pixel matrix $\mathbf{V}$ Gaussian matrix
$\mathbf{B}$ Exponent matrix $p(\mathcal{T})$ Center of tile $\mathcal{T}$
$\Delta p=(\Delta p_{x},\Delta p_{y}).$ Pixel on local coordinates $\Delta\mu^{\prime}=(\Delta\mu_{x}^{\prime},\Delta\mu_{y}^{\prime})$ Projected mean on local coordinates
$\varepsilon_{H}$ Machine epsilon of FP16 or TF32 $\varepsilon_{F}$ Machine epsilon of FP32
$R_{H}$ Rounding function of FP16 or TF32 $R_{F}$ Rounding function of FP32
$\hat{\beta}$ Computed value of $\beta$ with Tensor Cores $\lambda_{1},\lambda_{2}$ Two eigenvalues of $\Sigma^{\prime}$
Table 8. Notations in this paper.